OpenStudy (calculusxy):

An amateur rocketry club is holding a competition. There is a cloud cover at 1000 ft. If a rocket is launched with a velocity of 315ft/s, use the function h(t) = -16t^2+ vt + h_0 to determine how long the rocket is out of sight.

OpenStudy (calculusxy):

@phi @jim_thompson5910

jimthompson5910 (jim_thompson5910):



OpenStudy (calculusxy):

-16t^2 + 315t + 1000 = 0

jimthompson5910 (jim_thompson5910):

imagine the rocket taking off and following this path 

jimthompson5910 (jim_thompson5910):

the height of the rocket at time t is h(t) = -16t^2+ vt + h_0 h(t) = -16t^2+ 315t + 0 h(t) = -16t^2+ 315t set this height equal to 1000 h(t) = -16t^2+ 315t 1000 = -16t^2+ 315t now solve for t